The length of a rectangle is 8 inches less than 4 times the width. The perimeter is 114 inches. Find the length and the width.

The length is______ inches and the width is ______ inches

Answer by lwsshak3(11628)

The length of a rectangle is 8 inches less than 4 times the width. The perimeter is 114 inches. Find the length and the width.
***
let x=width
4x-8=length
2(length+width)=perimeter
2(x+4x-8)=114
10x-16=114
10x=130
x=13
4x-8=44
width=13 inches
length=44 inches